Shelhorse Corporation produces and sells a single product. Data concerning that product appear below: Per Unit Percent of Sales Selling price $ 240 100 % Variable expenses 120 50 % Contribution margin $ 120 50 % Fixed expenses are $352,000 per month. The company is currently selling 4700 units per month. 


Required: 

Calculate the Following: Contribution Margin Ratio: % Break Even point in Sales Dollars (rounded to the whole dollar) $ Margin of Safety in Units (nearest whole units) Units
Answer #1

Contribution margin ratio = 50%

Breakeven point in sales dollars

= Fixed cost/Contribution margin ratio

= 352,000/50%

= 704,000

Breakeven units = 704,000/240 = 2933 units

Margin of Safety = Sales - Breakeven sales

= 4700 - 2933

= 1767 units